OpenStudy (anonymous):

Please help :) What are the foci of the hyperbola given by the equation 16x^2 - 25y^2 - 64x - 50y - 361 = 0?

OpenStudy (anonymous):

(4(x-2))^2+(5(y-1))^2=9^2

OpenStudy (anonymous):

(4(x-2))^2+(5(y-1))^2=81

OpenStudy (anonymous):

x=2 y=1 r=9

OpenStudy (anonymous):

Um that's not a choice....